OpenStudy (anonymous):

1-cot^2 x / 1+cot^2 x = 1-2cos^2x how do I prove this ?

myininaya (myininaya):

\[\frac{1-\cot^2(x)}{1+\cot^2(x)}=\frac{1-\frac{\cos^2(x)}{\sin^2(x)}}{1+\frac{\cos^2(x)}{\sin^2(x)}}\] \[\frac{1-\frac{\cos^2(x)}{\sin^2(x)}}{1+\frac{\cos^2(x)}{\sin^2(x)}} \cdot \frac{\sin^2(x)}{\sin^2(x)}\] \[\frac{\sin^2(x)-\cos^2(x)}{\sin^2(x)+\cos^2(x)}=\frac{\sin^2(x)-\cos^2(x)}{1}=\sin^2(x)-\cos^2(x)\] \[(1-\cos^2(x))-\cos^2(x)=1-\cos^2(x)-\cos^2(x)=1-2\cos^2(x)\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

sin^2(x)/sin2^(x) why did you multiply this ?